# Approximability of Deep Computations

### v3

@@ 3.5.1. The unit interval
Let Γ¯=($\underline{\Gamma}=(\Gamma,\circ)$ be any semigroup (under composition) of continuous functions $\gamma:[0,1]\to[0,1]$, acting on $[0,1]$ by functional application: $\operatorname{ev}(\gamma,v)\coloneqq\gamma(v)$; this yields a realcompact CCS $\langle([0,1],P_{\operatorname{id}}),\underline{\Gamma}\rangle$. An interesting such CCS has semigroup $\Gamma=\{\gamma^{n}:n\in\mathbb{N}\}$ consisting of iterates of the chaotic map $\gamma:v\mapsto 4v(1-v)$.
@@ 3.5.2. $\mathbb{R}^{d}$
  • Γ¯=($\overline{\Gamma}=(\Gamma,\circ)$ any semigroup of continuous functions on $\mathbb{R}^{d}$; and
  • the evaluation action of $\Gamma$ on $L$ by functional application $\operatorname{ev}(\gamma,v)\coloneqq\gamma(v)$ as in 3.5.1 above,
@@ 3.5.3. $\mathbb{R}^{\omega}$ and $c_{00}$
Note that in the CSSs 3.5.13.5.3 above (but not in 3.5.4 below), a state $v$ is exactly the same as its type $\operatorname{tp}(v)$.
@@ 3.5.4. $\ell_{q}$
For $q<\infty$, such space $\ell_{q}$ is the $\left\|\cdot\right\|_{q}$-metric completion of the subspace $c_{00}\subseteq\mathbb{R}^{\omega}$; at any rate, $\ell_{\infty}$ is $\left\|\cdot\right\|_{\infty}$-complete as well. [^18] The $\left\|\cdot\right\|_{\infty}$-metric completion of $c_{00}$ is the separable space $c_{0}=\{v\in\mathbb{R}^{\omega}:\lim_{i\to\infty}v_{i}=0\}\subsetneq\ell_{\infty}$ —the space $\ell_{\infty}$ is not separable. A natural collection of predicates is $\mathcal{P}=(N_{q},P_{i})_{i\in\mathbb{N}}$, where $P_{i}:v\mapsto v_{i}$ is the $i$-th coordinate as in 3.5.3 above, and $N_{q}$ names the norm $N_{q}(\cdot):v\mapsto\left\|v\right\|_{q}$. Since the predicate collection $\mathcal{P}$ is countable, it is easy to show that the type topology and the usual $\left\|\cdot\right\|_{q}$-norm topology on the layer space $\ell_{q}$ coincide; [^19] Cf., the proof of Proposition 4.1 below. however, $\ell_{q}$ is non-realcompact. It is easy to see that its type space is $\mathfrak{L}_{q}=\{(v,r)\in\ell_{q}\times\mathbb{R}:r\geq\left\|v\right\|_{q}\}$. (The set of realized types is $\operatorname{tp}(\ell_{q})=\{(v,r)\in\ell_{q}\times\mathbb{R}:r=\left\|v\right\|_{q}\}\subseteq\mathfrak{L}_{q}$, for which the “correct” norm $\left\|v\right\|_{q}$ agrees with the interpretation value $N_{q}(v,r)=r$ of the symbol $N_{q}$.) Fixing $q<\infty$, the function $\left\|\cdot\right\|_{\infty}:\ell_{q}\to\mathbb{R}:v\mapsto\sup_{n}\left|P_{n}(v)\right|$ is $1$-Lipschitz, hence continuous on $\ell_{q}$; however, the corresponding function $\operatorname{tp}:\ell_{q}\to\mathbb{R}:\operatorname{tp}(v)\mapsto\left\|v\right\|_{\infty}$ does not extend continuously to $\mathfrak{L}_{q}$. [^20] By the Stone-Weierstrass Theorem, every continuous function on the compact Hausdorff space $\mathfrak{L}_{q}[1]\coloneqq\{(v,r)\in\mathfrak{L}_{q}:r\leq 1\}$ is uniformly approximable by algebraic combinations (finitely many at a time) of predicates $P_{i}$, and $\left\|\cdot\right\|_{q}$; however, an elementary argument shows that $\operatorname{tp}(v)\mapsto\left\|v\right\|_{\infty}$ admits no such uniform approximations on $\mathfrak{L}_{q}[1]$.
@@ 4.1.2. Shards
    If $S\subseteq\mathfrak{L}$ ($=\mathfrak{L}_{\mathrm{Sh}}$) is not closed, let $\mathfrak{u}\in\overline{S}\setminus S\subseteq\mathfrak{L}\setminus S$. As in (1) above, construct $(\mathfrak{v}_{n})_{n\in\mathbb{N}}\subseteq S$ with $\mathfrak{u}=\lim_{n\to\infty}(\mathfrak{v}_{n})$. Since $\overline{L}=\mathfrak{L}\supseteq S$, for each $n,k\in\mathbb{N}$ there exists some $v^{(n)}_{k}\in L$ such that $\operatorname{d}(v^{(n)}_{k},\mathfrak{v}_{n})<1/(n+k)$. One sees that $A=\{\operatorname{tp}(v^{(n)}_{k}):n,k\in\mathbb{N}\}\cup\{\mathfrak{v}_{n}:n\in\mathbb{N}\}\cup\{\mathfrak{u}\}$ is compact. [^22] Clearly, $A^{(n)}\coloneqq\{\mathfrak{v}_{n}\}\cup\{\operatorname{tp}(v^{(n)}_{k}):k\in\mathbb{N}\}$ is compact for each $n\in\mathbb{N}$ (as in (1) above). Given any open cover $\mathcal{O}$ of $A$, some $U\in\mathcal{O}$ satisfies $U\ni\mathfrak{u}$. For a set $N\subseteq\mathbb{N}$ containing all but finitely many $n$, we have $U\supseteq A^{(n)}$. The set $B\coloneqq\bigcup_{n\in\mathbb{N}\setminus N}A^{(n)}$ is a finite union of compacta, hence itself compact. Thus, finitely many opens of $\mathcal{O}$ cover $B$, which together with $U$ cover $A$. Therefore, $A\subseteq\mathfrak{L}[r_{\bullet}]$ for some sizer $r_{\bullet}$, and thus $\{v^{(n)}_{k}:n,k\in\mathbb{N}\}\subseteq L[r_{\bullet}]$, hence $\mathfrak{u}=\lim_{n\to\infty}\operatorname{tp}(v^{(n)}_{n})\in\overline{\operatorname{tp}(L[r_{\bullet}])}=\mathfrak{L}[r_{\bullet}]$. Since $\mathfrak{u}\in\overline{S}\setminus S$ and $S\subseteq\mathfrak{L}[r_{\bullet}]$ we see that $S\cap\mathfrak{L}[r_{\bullet}]$ is not closed. The converse is trivial, since type-shards $\mathfrak{L}[r_{\bullet}]$ are closed (and the intersection of closed sets is closed). [^23] We recall the following fact, closely related to (2): Every sequential Hausdorff (therefore, every metric) space is a k-space [Eng89, Theorem 3.3.20].
@@ 5.1. Layered and iterative computations
The term “evolution” means “$\gamma_{\bullet}$-evolution” henceforth, whenever $\gamma_{\bullet}$ is given by context. The state at stage $n$ of $v$ under evolution is $\operatorname{ev}(\gamma^{(n)},v)$.
@@ 5.3. Examples and discussion of deep iterates and deep equilibria
$\begin{split}f^{*}(f^{*}(i))&=f^{N}(f^{N-n}(f^{n}(i)))=g^{N}(g^{N-n}(f^{n}(i)))\\ &\qquad\text{(since $f^{n}([m])\subseteq[k]$ and $g=f\restriction[k]$)}\\ &=g^{N-n}(g^{N}(f^{n}(i)))=g^{N-n}(f^{n}(i))\\ &\qquad\text{(since $K$ divides $N$ and $g^{K}$ is the identity)}\\ &=f^{N-n}(f^{n}(i))=f^{N}(i)=f^{*}(i).\end{split}$
@@
(^,$\bigl{\langle}(\hat{\mathbb{C}},\{U,V,W\}),\langle f\rangle\bigr{\rangle},$
@@
  • ${}^\hat{\mathbb{C}}=\mathbb{C}\cup\{\infty\}$ is the Riemann sphere, which we identify with the unit sphere $S^{2}=\{(u,v,w)\in\mathbb{R}^{3}:u^{2}+v^{2}+w^{2}=1\}$ via, e.g., the stereographic projection $(u,v,w)\mapsto z=(u+iv)/(1-w)$ (and $(0,0,1)\mapsto\infty$);
  • ^$\hat{\mathbb{C}}\to S^{3}:z\mapsto(U,V,W)$ is the inverse of the stereographic projection, regarded as a triple of predicates $U,V,W:\hat{\mathbb{C}}\to[-1,1]$; and
@@
Since $\hat{\mathbb{C}}$ is compact and $z\mapsto(U(z),V(z),W(z))$ is a homeomorphic embedding, In fact, $\hat{\mathbb{C}}$ is equal to the shard ^$\hat{\mathbb{C}}[1,1,1]=\bigl{\{}z:\max\bigl{(}\left|U(z)\right|,\left|V(z)\right|,\left|W(z)\right|\bigr{)}\leq 1\bigr{\}}$ (in particular, $\hat{\mathbb{C}}$ is realcompact); thus, $\gamma$ is automatically confined (by $R$ consisting of the single sizer $r_{\bullet}=(r_{U},r_{V},r_{W})=(1,1,1)$).
Let $\gamma^{*}$ be any deep iterate of $\gamma$. At any point $z\in\hat{\mathbb{C}}$ for which the Newton method converges to a root $w$ of $p(z)$ (in particular, at any $z$ sufficiently close to a simple such root $w$), we have $\gamma^{*}(z)=w$ ($=\gamma(w)$, since $p(w)=0$). We also have $\gamma^{*}(\infty)=\infty=\gamma(\infty)$; however, $\infty$ is a repeller [^28] Perhaps surprisingly, it is possible for the fixed repeller $\infty$ to be an accumulation point of orbits $(\gamma^{n}(z))_{n\in\mathbb{N}}\subseteq\mathbb{C}$. This is the case, e.g., for the polynomial $p(z)=z^{3}-1$. (this follows from the easy calculation that $p(z)/z\to 1-d^{-1}$ as $\left|z\right|\to+\infty$), so one would expect points $z\in\mathbb{C}$ with $\gamma^{*}(z)=\infty$ to be quite scant. In general, however, $w\coloneqq\gamma^{*}(z)$ is not a root of $p$, although any such $w\in\hat{\mathbb{C}}$ is necessarily a topologically recurrent point of $\hat{\mathbb{C}}$ under $\gamma$. At any rate, if $p$ has at least two distinct roots, any deep equilibrium (or deep iterate) $\gamma^{*}$ of $\gamma$ is discontinuous.
@@
    Even in situations where, say, a deep iterate does not quite exist, an ultracomputation may have “meaningful deep features” in a sense that we now explain. Consider any CCS $\mathcal{C}=(L,\langle f\rangle,\circ,\mathcal{P})$ (not necessarily realcompact), where $f:L\to L$ is any given (continuous) computation. For a fixed $Q\in\mathcal{P}$, say that $f$ has uniformly $Q$-bounded iterates on $v\in L$ if there exists $s=s^{(v)}>0$ such that $\left|Q(f^{n}(v))\right|\leq s$ for all $n\in\mathbb{N}$. (Note that this hypothesis does not—at all—impose bounds on other entries $P\circ f^{n}(v)$ for $Q\neq P\in\mathcal{P}$.) If $\mathcal{U}$ is any nonprincipal ultrafilter on $\mathbb{N}$, the iterate boundedness hypothesis and the compactness of intervals $[-s,s]$ imply that $\operatorname{\mathcal{U}lim}Q(f^{n}(v))$ exists for all $v\in L$. In principle, however, the iterates $f^{n}$ need not $\mathcal{U}$-converge in $T$ (i.e., pointwise on $L$) even if $L$ is realcompact, since (for fixed $v$) the sequence $(f^{n}(v))_{n\in\mathbb{N}}$ may not be entry-wise bounded (only bounded “in $Q$-th entry”, so to speak).
@@ 6. Explicit computability
    • Addition:$(\varphi,\psi)\mapsto\varphi+\psi$ (where $\varphi+\psi:v\mapsto\varphi(v)+\psi(v)$);
    • Multiplication:$(\varphi,\psi)\mapsto\varphi\psi$ (where $\varphi\psi:v\mapsto\varphi(v)\psi(v)$).
@@
    By the definition of the topologies on $L$ and $\mathfrak{L}_{\mathrm{Sh}}$, every monomial $P$ is continuous and bounded by $r_{P}$ on any shard $L[r_{\bullet}]$, and extends continuously to $\mathfrak{L}_{\mathrm{Sh}}$ (as the $P$-coordinate function). Since connectives are obtained by pointwise application of continuous real-valued functions of real arguments (addition and multiplication), every polynomial on $L$ is also continuous, and extends to a continuous bounded function on type-shards $\mathfrak{L}[r_{\bullet}]$. Definable predicates, on the other hand, need not be continuous on $L$—although their restrictions to shards $L[r_{\bullet}]$ necessarily are continuous and bounded (being uniform limits of polynomial on $L[r_{\bullet}]$, which is compact).
@@
    Let $r_{\bullet}$ be an arbitrary sizer. The restriction of a monomial $P$ to the type-shard $\mathfrak{L}[r_{\bullet}]$ admits the a priori bound $C=r_{P}$, so that that $P{\restriction}\mathfrak{L}[r_{\bullet}]$ takes values in $[-C,C]=[-r_{P},r_{P}]$. [^32] A constant $r$ also admits the trivial bound $C=\left|r\right|$. By recursion on the application of connectives leading to an arbitrary polynomial $\varphi$ from monomials, a priori bounds $C=C^{\varphi}_{r_{\bullet}}\in[0,\infty)$ such that $\varphi{\restriction}\mathfrak{L}[r_{\bullet}]$ takes values in $[-C,C]$ are easily found. (Recursively apply the rules: $C_{r_{\bullet}}^{\varphi+\psi}\coloneqq C_{r_{\bullet}}^{\varphi}+C_{r_{\bullet}}^{\psi}$, and $C_{r_{\bullet}}^{\varphi\psi}\coloneqq C_{r_{\bullet}}^{\varphi}\cdot C_{r_{\bullet}}^{\psi}$.)

### v5

@@ 2.5.1. The unit interval
Let $\underline{\Gamma}=(\Gamma,\circ)$ be any semigroup (under composition) of continuous functions $\gamma:[0,1]\to[0,1]$, acting on $[0,1]$ by functional application: $\operatorname{ev}(\gamma,v)\coloneqq\gamma(v)$; this yields a realcompact CCS $\langle([0,1],P_{\operatorname{id}}),\underline{\Gamma}\rangle$. An interesting such CCS has semigroup $\Gamma=\{\gamma^{n}:n\in\mathbb{N}\}$ consisting of iterates of the chaotic map $\gamma:v\mapsto 4v(1-v)$.
@@ 2.5.2. $\mathbb{R}^{d}$
  • $\overline{\Gamma}=(\Gamma,\circ)$ any semigroup of continuous functions on $\mathbb{R}^{d}$; and
  • the evaluation action of $\Gamma$ on $L$ by functional application $\operatorname{ev}(\gamma,v)\coloneqq\gamma(v)$ as in 2.5.1 above,
@@ 2.5.3. $\mathbb{R}^{\omega}$ and $c_{00}$
Note that in the CSSs 2.5.12.5.3 above (but not in 2.5.4 below), a state $v$ is exactly the same as its type $\operatorname{tp}(v)$.
@@ 2.5.4. $\ell_{q}$
For $q<\infty$, such space $\ell_{q}$ is the $\left\|\cdot\right\|_{q}$-metric completion of the subspace $c_{00}\subseteq\mathbb{R}^{\omega}$; at any rate, $\ell_{\infty}$ is $\left\|\cdot\right\|_{\infty}$-complete as well. [^16] The $\left\|\cdot\right\|_{\infty}$-metric completion of $c_{00}$ is the separable space $c_{0}=\{v\in\mathbb{R}^{\omega}:\lim_{i\to\infty}v_{i}=0\}\subsetneq\ell_{\infty}$ —the space $\ell_{\infty}$ is not separable. A natural collection of predicates is $\mathcal{P}=(N_{q},P_{i})_{i\in\mathbb{N}}$, where $P_{i}:v\mapsto v_{i}$ is the $i$-th coordinate as in 2.5.3 above, and $N_{q}$ names the norm $N_{q}(\cdot):v\mapsto\left\|v\right\|_{q}$. Since the predicate collection $\mathcal{P}$ is countable, it is easy to show that the type topology and the usual $\left\|\cdot\right\|_{q}$-norm topology on the layer space $\ell_{q}$ coincide; [^17] Cf., the proof of Proposition 3.1 below. however, $\ell_{q}$ is non-realcompact. It is easy to see that its type space is $\mathfrak{L}_{q}=\{(v,r)\in\ell_{q}\times\mathbb{R}:r\geq\left\|v\right\|_{q}\}$. (The set of realized types is $\operatorname{tp}(\ell_{q})=\{(v,r)\in\ell_{q}\times\mathbb{R}:r=\left\|v\right\|_{q}\}\subseteq\mathfrak{L}_{q}$, for which the “correct” norm $\left\|v\right\|_{q}$ agrees with the interpretation value $N_{q}(v,r)=r$ of the symbol $N_{q}$.) Fixing $q<\infty$, the function $\left\|\cdot\right\|_{\infty}:\ell_{q}\to\mathbb{R}:v\mapsto\sup_{n}\left|P_{n}(v)\right|$ is $1$-Lipschitz, hence continuous on $\ell_{q}$; however, the corresponding function $\operatorname{tp}:\ell_{q}\to\mathbb{R}:\operatorname{tp}(v)\mapsto\left\|v\right\|_{\infty}$ does not extend continuously to $\mathfrak{L}_{q}$. [^18] By the Stone-Weierstrass Theorem, every continuous function on the compact Hausdorff space $\mathfrak{L}_{q}[1]\coloneqq\{(v,r)\in\mathfrak{L}_{q}:r\leq 1\}$ is uniformly approximable by algebraic combinations (finitely many at a time) of predicates $P_{i}$, and $\left\|\cdot\right\|_{q}$; however, an elementary argument shows that $\operatorname{tp}(v)\mapsto\left\|v\right\|_{\infty}$ admits no such uniform approximations on $\mathfrak{L}_{q}[1]$.
@@ 3.1.2. Shards
    If $S\subseteq\mathfrak{L}$ ($\mathfrak{L}_{\mathrm{Sh}}$) is not closed, let $\mathfrak{u}\in\overline{S}\setminus S\subseteq\mathfrak{L}\setminus S$. As in (1) above, construct $(\mathfrak{v}_{n})_{n\in\mathbb{N}}\subseteq S$ with $\mathfrak{u}=\lim_{n\to\infty}(\mathfrak{v}_{n})$. Since $\overline{L}=\mathfrak{L}\supseteq S$, for each $n,k\in\mathbb{N}$ there exists some $v^{(n)}_{k}\in L$ such that $\operatorname{d}(v^{(n)}_{k},\mathfrak{v}_{n})<1/(n+k)$. One sees that $A=\{\operatorname{tp}(v^{(n)}_{k}):n,k\in\mathbb{N}\}\cup\{\mathfrak{v}_{n}:n\in\mathbb{N}\}\cup\{\mathfrak{u}\}$ is compact. [^20] Clearly, $A^{(n)}\coloneqq\{\mathfrak{v}_{n}\}\cup\{\operatorname{tp}(v^{(n)}_{k}):k\in\mathbb{N}\}$ is compact for each $n\in\mathbb{N}$ (as in (1) above). Given any open cover $\mathcal{O}$ of $A$, some $U\in\mathcal{O}$ satisfies $U\ni\mathfrak{u}$. For a set $N\subseteq\mathbb{N}$ containing all but finitely many $n$, we have $U\supseteq A^{(n)}$. The set $B\coloneqq\bigcup_{n\in\mathbb{N}\setminus N}A^{(n)}$ is a finite union of compacta, hence itself compact. Thus, finitely many opens of $\mathcal{O}$ cover $B$, which together with $U$ cover $A$. Therefore, $A\subseteq\mathfrak{L}[r_{\bullet}]$ for some sizer $r_{\bullet}$, and thus $\{v^{(n)}_{k}:n,k\in\mathbb{N}\}\subseteq L[r_{\bullet}]$, hence $\mathfrak{u}=\lim_{n\to\infty}\operatorname{tp}(v^{(n)}_{n})\in\overline{\operatorname{tp}(L[r_{\bullet}])}=\mathfrak{L}[r_{\bullet}]$. Since $\mathfrak{u}\in\overline{S}\setminus S$ and $S\subseteq\mathfrak{L}[r_{\bullet}]$ we see that $S\cap\mathfrak{L}[r_{\bullet}]$ is not closed. The converse is trivial, since type-shards $\mathfrak{L}[r_{\bullet}]$ are closed (and the intersection of closed sets is closed). [^21] We recall the following fact, closely related to (2): Every sequential Hausdorff (therefore, every metric) space is a k-space [Eng89, Theorem 3.3.20].
@@ 4.1. Layered and iterative computations
The term “evolution” means $\gamma_{\bullet}$-evolution” henceforth, whenever $\gamma_{\bullet}$ is given by context. The state at stage $n$ of $v$ under evolution is $\operatorname{ev}(\gamma^{(n)},v)$.
@@ 4.3. Examples and discussion of deep iterates and deep equilibria
[…]
@@
$\bigl\langle(\hat{\mathbb{C}},\{U,V,W\}),\langle f\rangle\bigr\rangle,$
@@
  • $\hat{\mathbb{C}}$$\mathbb{C}\cup\{\infty\}$ is the Riemann sphere, which we identify with the unit sphere $S^{2}=\{(u,v,w)\in\mathbb{R}^{3}:u^{2}+v^{2}+w^{2}=1\}$ via, e.g., the stereographic projection $(u,v,w)\mapsto z=(u+iv)/(1-w)$ (and $(0,0,1)\mapsto\infty$);
  • $\hat{\mathbb{C}}\to S^{3}:z\mapsto(U,V,W)$ is the inverse of the stereographic projection, regarded as a triple of predicates $U,V,W:\hat{\mathbb{C}}\to[-1,1]$; and
@@
Since $\hat{\mathbb{C}}$ is compact and $z\mapsto(U(z),V(z),W(z))$ is a homeomorphic embedding, In fact, $\hat{\mathbb{C}}$ is equal to the shard $\hat{\mathbb{C}}[1,1,1]=\bigl\{z:\max\bigl(\left|U(z)\right|,\left|V(z)\right|,\left|W(z)\right|\bigr)\leq 1\bigr\}$ (in particular, $\hat{\mathbb{C}}$ is realcompact); thus, $\gamma$ is automatically confined (by $R$ consisting of the single sizer $r_{\bullet}=(r_{U},r_{V},r_{W})=(1,1,1)$).
Let $\gamma^{*}$ be any deep iterate of $\gamma$. At any point $z\in\hat{\mathbb{C}}$ for which the Newton method converges to a root $w$ of $p(z)$ (in particular, at any $z$ sufficiently close to a simple such root $w$), we have $\gamma^{*}(z)=w$ ($\gamma(w)$, since $p(w)=0$). We also have $\gamma^{*}(\infty)=\infty=\gamma(\infty)$; however, $\infty$ is a repeller [^26] Perhaps surprisingly, it is possible for the fixed repeller $\infty$ to be an accumulation point of orbits $(\gamma^{n}(z))_{n\in\mathbb{N}}\subseteq\mathbb{C}$. This is the case, e.g., for the polynomial $p(z)=z^{3}-1$. (this follows from the easy calculation that $p(z)/z\to 1-d^{-1}$ as $\left|z\right|\to+\infty$), so one would expect points $z\in\mathbb{C}$ with $\gamma^{*}(z)=\infty$ to be quite scant. In general, however, $w\coloneqq\gamma^{*}(z)$ is not a root of $p$, although any such $w\in\hat{\mathbb{C}}$ is necessarily a topologically recurrent point of $\hat{\mathbb{C}}$ under $\gamma$. At any rate, if $p$ has at least two distinct roots, any deep equilibrium (or deep iterate) $\gamma^{*}$ of $\gamma$ is discontinuous.
@@
    Even in situations where, say, a deep iterate does not quite exist, an ultracomputation may have “meaningful deep features” in a sense that we now explain. Consider any CCS $\mathcal{C}=(L,\langle f\rangle,\circ,\mathcal{P})$ (not necessarily realcompact), where $f:L\to L$ is any given (continuous) computation. For a fixed $Q\in\mathcal{P}$, say that $f$ has uniformly $Q$-bounded iterates on $v\in L$ if there exists $s=s^{(v)}>0$ such that $\left|Q(f^{n}(v))\right|\leq s$ for all $n\in\mathbb{N}$. (Note that this hypothesis does not—at all—impose bounds on other entries $P\circ f^{n}(v)$ for $Q\neq P\in\mathcal{P}$.) If $\mathcal{U}$ is any nonprincipal ultrafilter on $\mathbb{N}$, the iterate boundedness hypothesis and the compactness of intervals [s,s][-s,s] imply that $\operatorname{\mathcal{U}lim}Q(f^{n}(v))$ exists for all $v\in L$. In principle, however, the iterates $f^{n}$ need not $\mathcal{U}$-converge in $T$ (i.e., pointwise on $L$) even if $L$ is realcompact, since (for fixed $v$) the sequence $(f^{n}(v))_{n\in\mathbb{N}}$ may not be entry-wise bounded (only bounded “in $Q$-th entry”, so to speak).
@@ 5. Explicit computability
[…]
@@
    By the definition of the topologies on $L$ and $\mathfrak{L}_{\mathrm{Sh}}$, every monomial $P$ is continuous and bounded by $r_{P}$ on any shard $L[r_{\bullet}]$, and extends continuously to $\mathfrak{L}_{\mathrm{Sh}}$ (as the $P$-coordinate function). Since connectives are obtained by pointwise application of continuous real-valued functions of real arguments (addition and multiplication), every polynomial on $L$ is also continuous, and extends to a continuous bounded function on type-shards $\mathfrak{L}[r_{\bullet}]$. Definable predicates, on the other hand, need not be continuous on LL—although their restrictions to shards $L[r_{\bullet}]$ necessarily are continuous and bounded (being uniform limits of polynomial on $L[r_{\bullet}]$, which is compact).
@@
    Let $r_{\bullet}$ be an arbitrary sizer. The restriction of a monomial $P$ to the type-shard $\mathfrak{L}[r_{\bullet}]$ admits the a priori bound $C=r_{P}$, so that that $P{\restriction}\mathfrak{L}[r_{\bullet}]$ takes values in $C,C]=[-r_{P},r_{P}]$. [^30] A constant $r$ also admits the trivial bound $C=\left|r\right|$. By recursion on the application of connectives leading to an arbitrary polynomial $\varphi$ from monomials, a priori bounds $C=C^{\varphi}_{r_{\bullet}}\in[0,\infty)$ such that $\varphi{\restriction}\mathfrak{L}[r_{\bullet}]$ takes values in [C,C][-C,C] are easily found. (Recursively apply the rules: $C_{r_{\bullet}}^{\varphi+\psi}\coloneqq C_{r_{\bullet}}^{\varphi}+C_{r_{\bullet}}^{\psi}$, and $C_{r_{\bullet}}^{\varphi\psi}\coloneqq C_{r_{\bullet}}^{\varphi}\cdot C_{r_{\bullet}}^{\psi}$.)